OpenStudy (anonymous):

jenny writes the sequence 3 7 11 15...

OpenStudy (anonymous):

just keep adding 4 to the next numbers 3+4=7+4=11+4=15+4=19+4=23 see what I am getting at?

OpenStudy (texaschic101):

good for jenny....now what is the question ?